Geoffrey C. Fox

Phone: +1 (315) 443 2163 Fax: +1 (315) 443 1973 E-mail: gcf@npac.syr.edu

August 29, 1997

Rebecca A. McKinley Subcontracts Supervisor Nichols Research Corporation 4040 South Memorial Parkway Huntsville, AL 35802

Dear Ms. McKinley:

After discussions with Dick Pritchard, PET Director at CEWES, we would like to purchase the computer equipment described below on subcontract NRC-CR-96-0005 (DoD/HPC Modernization Program at the CEWES MSRC, Vicksburg, Mississippi). This equipment will facilitate our work on all aspects of this project, but most especially it will benefit our “Focused Effort on Distance Education for Curricula Enhancement”, which is of great interest to CEWES.

The proposed equipment consists of four SGI workstations and two PC systems, fully equipped for multimedia operation. The attached price quotes provide both more detailed information about the equipment configurations and the costs. Please note that purchasing additional memory for the O2 workstations from a third party results in savings of more than $2000. The dollar amounts are listed in the table below, and of course in the quotes. The costs of this equipment will be paid by reallocating current salary funding on our subcontract – no additional funding is required for this purchase. This reallocation will improve out ability to deliver according to the Statement of Work in our subcontract.

Vendor  Description  Unit Price Qty  Total
Silicon Graphics O2 workstations
w/ 64 MB RAM, 4 GB HDD  2
w/ 64 MB RAM, 4 GB + 2 GB HDD  2
SGI Total  28177.60
Impediment Inc. 64 MB upgrade kit for O2  459.00  4 1836.00
Gateway2000  GP6-266 w/ multimedia capabilities  3794.00  2 7588.00
TOTAL  37601.60

In order to get this equipment in place as soon as posible, in support of a distance education course which is already underway, we would appreciate your prompt attention to this request.

Sincerely,

Geoffrey C. Fox

Director, NPAC and Professor of Physics and Computer Science